OpenStudy (anonymous):

please help solve: 4^(10x-10)=5^(5x-7)

OpenStudy (anonymous):

start with \[(10x-10)\ln(4)=(5x-7)\ln(5)\] and then do a bunch of algebra

myininaya (myininaya):

\[\ln(4^{10x-10})=\ln(5^{5x-7})\]

OpenStudy (anonymous):

last night: integrals out the yin yang tonight : algebra for days

OpenStudy (anonymous):

don't forget that whatever those numbers are, \[\ln(4)\text{ and } \ln(5)\] are just constants

myininaya (myininaya):

\[\ln(4) 10x-10 \ln(4)=\ln(5) 5x-7\ln(5)\]

OpenStudy (anonymous):

what

myininaya (myininaya):

put like terms together put x terms on one side and put constant terms on the other side

OpenStudy (anonymous):

\[\10\ln(4)x-5\ln(5)x=10\ln(4)-7\ln(5)\]

myininaya (myininaya):

now we can write left hand side as....

OpenStudy (anonymous):

and finally \[x(10\ln(4)-5\ln(5))=10\ln(4)-7\ln(5)\] divide to get ...

myininaya (myininaya):

and then \[x=\frac{10 \ln(4)-7 \ln(5)}{10\ln(4)-5\ln(5)}\]

OpenStudy (anonymous):

lol

myininaya (myininaya):

lol

myininaya (myininaya):

goodnight

OpenStudy (anonymous):

gnight

OpenStudy (anonymous):

thanx again both of you =)